The difference of two numbers is 91 The second is 5 less than 5 times the first What are the two numbers
Answer by checkley71(8403)

X-Y=91
X=91+Y AND
Y=5X-5 THEN
X=91+(5X-5)
X=91+5X-5
X-5X=86
-4X=86
X=86/-4
X=-21.5 SOLUTION
-21.5+Y=91
Y=91+12.5
Y=103.5 SOLUTION
PROOF
-21.5=91+(5*-21.5-5)
-21.5=91-107.5-5
-21.5=-21.5